# ObjectClear: Complete Object Removal via Object-Effect Attention

### v1

@@ Abstract
Object removal requires eliminating not only the target object but also its effects, such as shadows and reflections. However, diffusion-based inpainting methods often produce artifacts, hallucinate content, alter background, and struggle to remove object effects accurately. To address this challenge, we introduce a new dataset for OBject-Effect Removal, named OBER, which provides paired images with and without object effects, along with precise masks for both objects and their associated visual artifacts. The dataset comprises high-quality captured and simulated data, covering diverse object categories and complex multi-object scenes. Building on OBER, we propose a novel framework, ObjectClear, which incorporates an object-effect attention mechanism to guide the model toward the foreground removal regions by learning attention masks, effectively decoupling foreground removal from background reconstruction. Furthermore, the predicted attention map enables an attention-guided fusion strategy during inference, greatly preserving background details. Extensive experiments demonstrate that ObjectClear outperforms existing methods, achieving improved object-effect removal quality and background fidelity, especially in complex scenarios.
footnotetext: Corresponding author.
@@ 1 Introduction
Figure 1: Object Removal Comparison. Given an object mask, prior methods often leave residual artifacts or hallucinate undesirable content, change background, and typically fail to remove associated effects such as shadows and reflections. In contrast, our ObjectClear precisely eliminates both the object and its associated effects, achieving seamless object removal results even in challenging cases.
@@
Recent advances in generative models [27, 24, 21], such as GPT-4o, have shown strong capabilities in image editing and are widely adopted in real-world applications. Among these tasks, object removal has become a key task, allowing users to erase unwanted content from images. However, seamlessly removing both the object and its effects (e.g., shadows and reflections) while preserving the background remains a challenging problem.
While diffusion-based methods [1, 43, 12, 15, 31, 7, 10, 5] have advanced object removal, they still struggle with high-fidelity results. As shown in Fig. 1, existing methods often leave residual artifacts, hallucinate unwanted content, alter backgrounds details, and fail to remove object effects. This is partly because the input mask fully occludes the target object, preventing models from perceiving its surrounding context and associated effects. In addition, the lack of large-scale and public datasets for object-effect removal also limits further progress in the field.
@@
Existing object removal datasets can be categorized into simulated and camera-captured data. (1) Simulated data. These datasets are often constructed by copy-pasting objects [11, 14] or using pretrained inpainting models to generate pseudo ground truth [32]. While this allows low-cost generation of large-scale data, such datasets typically lack object effects such as shadows and reflections, causing models trained on them to struggle with effect removal. (2) Camera-captured data. Some works [29, 33] leverage fixed-viewpoint videos to extract paired frames, but this constrains the foreground types to moving objects. Others [34, 39, 36] capture image pairs before and after object removal, but these datasets are often not publicly available and small in scale due to the high costs.
@@
To bridge this gap, we introduce OBER, a hybrid dataset for OBject-Effect Removal, which combines both camera-captured and simulated data with diverse foreground objects, background scenes, and object effects. For the camera-captured data, we annotate precise object and object-effect masks, which serve as critical supervision during training. For the simulated data, we leverage these masks to compute accurate alpha maps, enabling realistic alpha blending of RGBA object layers with effects and high-quality backgrounds. We further extend the simulation pipeline to include multi-object scenarios, enhancing model robustness in challenging cases involving occlusion and object interactions, as shown in the last two rows of Fig. 5. By combining the realism of captured data with the scalability of simulation, OBER offers a high-quality, large-scale, and diverse dataset with a total of 12,715 training samples. In addition, we introduce two new benchmarks: OBER-Test (w/ ground truth) and OBER-Wild (w/o ground truth), to support future research.
Built upon OBER, we propose a novel framework, ObjectClear, which introduces an Object-Effect Attention (OEA) mechanism that adaptively focuses on the foreground removal region, including both the object and its associated effects, by supervising cross-attention maps with object-effect masks. The OEA module effectively decouples foreground removal from background reconstruction, enabling more precise and complete object elimination. In addition, the predicted attention map supports an Attention-Guided Fusion (AGF) strategy during inference, helping to preserve background details, as shown in Fig. 7. This design enables the model to localize editable regions adaptively and offers a potential solution to the long-standing fidelity issues in diffusion-based image editing.
Our contributions are summarized as follows:
  • We propose OBER, a high-quality and large-scale hybrid dataset for object removal, featuring diverse objects, fine-grained annotations of object-effect masks, and complex multi-object scenarios across both simulated and camera-captured settings.
  • We introduce a novel framework, ObjectClear, which incorporates an Object-Effect Attention mechanism that adaptively focuses on foreground removal regions and enables an Attention-Guided Fusion strategy, improving object removal quality and background fidelity.
  • We present two new benchmarks, OBER-Test and OBER-Wild, providing a foundation for future research. Our method achieves superior performance on all benchmarks, outperforming existing approaches in terms of both quantitative metrics and visual quality.
@@ 2 Related Work
Image Inpainting. Image inpainting is a long-standing visual editing task that aims to seamlessly reconstruct pixels within a given mask. Early approaches predominantly adopt generative adversarial networks (GANs) [8], but often suffer from limited realism and diversity [17, 22, 26, 40]. With the rapid advances in diffusion models [28, 30, 9, 24], many methods [2, 20, 21, 43, 12, 38] have begun leveraging their strong generative priors to synthesize high-fidelity content, achieving state-of-the-art results in image inpainting. In this work, we adapt the SDXL-Inpainting model [1] for photorealistic completion. However, despite their strong generative capabilities, existing inpainting models often lack awareness of object-induced effects (e.g., shadows and reflections), leading to incomplete or visually inconsistent object removal results.
Object Removal. Object removal is a specialized form of image inpainting that requires accounting for object effects to ensure complete and realistic removal, which diffusion-based models [34, 11, 19, 39, 15, 4, 43, 7, 5, 36, 10, 31, 33] currently dominate. A common strategy involves curating high-quality triplet datasets. For instance, ObjectDrop [34] builds a real-world dataset by having photographers capture the same scene before and after the removal of a single object. However, its limited scale and lack of public availability hinder broader adoption. To address data scalability, methods like SmartEraser [11] and Erase Diffusion [19] rely on synthetic datasets using segmentation or matting techniques to extract foreground objects, but these typically lack annotations of object effects, limiting the models’ ability to remove shadows or reflections. To enhance realism, methods such as LayerDecomp [36] and OmniPaint [39] create costly camera-captured data. OmniPaint [39] takes one step further to annotate unlabeled images by a model trained on small-scale real data, while RORem [15] involves human annotators to ensure the quality of annotated data. In parallel, works like RORD [29] and OmniEraser [33] scale data generation by mining realistic video frames with fixed viewpoints, selectively pairing frames with and without target objects while preserving natural object effects. To eliminate reliance on curated datasets altogether, other methods opt for test-time optimization [5, 31, 10] or providing plug-and-play solutions [7]. However, all these models rely on implicit learning of object effects without explicitly modeling effect maps, making it difficult to ensure that no unnecessary background changes occur.
@@ 3 Methodology
To achieve accurate removal of target objects along with their associated visual effects, such as shadow and reflection, we propose a comprehensive framework consisting of a data curation pipeline (OBER Dataset) and a dedicated object removal model (ObjectClear). These two components are detailed in Sec. 3.1 and Sec. 3.2, respectively.
@@ 3.1 OBER Dataset
ObjectDrop [34] shows that even small-scale datasets, when captured with careful camera control over scene consistency, can significantly enhance model generalizability in object removal tasks. However, due to the lack of public access to such datasets and the high cost of real-world data collection, subsequent works often face limitations in dataset scale. To address this challenge, we introduce the OBER dataset (OBject-Effect Removal), designed to balance data realism and scalability. As illustrated in Fig. 2, OBER is a hybrid dataset consisting of two parts: (1) a small set of camera-captured images adhering to physical realism, and (2) a larger set of simulated images generated by compositing foreground objects, extracted from the real data in (1), onto diverse background scenes.
Camera-Captured Data. (1) Capture Paired Images. Following the approach of ObjectDrop [34], we use fixed cameras to construct a counterfactual dataset by capturing each scene before and after the removal of a single object, while keeping all the other factors unchanged. For each pair, the image with the object is used as the input $I_{in}$, and the image without the object serves as the ground truth $I_{gt}$. In total, we collected 2,875 such counterfactual pairs with 2,715 for training and 160 for testing. Our dataset encompasses a wide variety of everyday objects commonly found in indoor (e.g., benches, cups, dolls) and outdoor scenes (e.g., pedestrians, vehicles). These image pairs also preserve visual effects such as shadows and reflections. To reduce potential pixel-level misalignment, we downsample and crop all images to a fixed training resolution (i.e., 512 $\times$ 512). (2) Segment Object Mask $M_{o}$. To obtain the object mask $M_{o}$ as network input alongside the input image $I_{in}$, we apply off-the-shelf detection and segmentation models, such as DINO [41] and SAM [13] to $I_{in}$. (3) Segment Object-Effect Mask $M_{o}\cup M_{e}$. Different from previous methods, we propose to explicitly model the object effects. Therefore, we introduce an object-effect mask $M_{fg}$ that covers both the object $M_{o}$ and its associated visual effects $M_{e}$ in our dataset. Unlike the coarse object-effect masks provided in the RORD dataset [29], we construct $M_{fg}$ in an efficient and accurate way by computing the pixel-wise difference between $I_{in}$ and $I_{gt}$. Pixels with differences above a predefined threshold are regarded as part of the object-effect mask. This object-effect mask provides crucial supervision during training, allowing the network to adaptively learn to focus on the target removal regions, including both the object itself and its effects.
@@
$\alpha(p)=\begin{cases}0,&\text{if pixel }p\in\text{background area $\overline{M_{o}}\cap\overline{M_{e}}$}\\ 1,&\text{if pixel }p\in\text{object area $M_{o}$}\\ (I_{\text{gt}}-I_{\text{in}})/(I_{\text{gt}}+\varepsilon),&\text{if pixel }p\in\text{effect area $M_{e}$}\\ \end{cases}$ (1)
(3) Blend Objects with Backgrounds. We randomly sample a background image $I_{bg}$ and a foreground object image with effects $I_{oe}$, and synthesize a composite image using alpha blending: $I_{comp}=(1-\alpha)\cdot I_{bg}+\alpha\cdot I_{oe}$. Beyond single-object cases, we also synthesize multi-object data by compositing multiple foreground objects with the same lighting directions, thereby covering scenarios involving object occlusions, as illustrated in Fig. 2(b). In total, we generate 10,000 composite images, significantly enriching the diversity and scalability of our dataset. In particular, the simulation of multi-object compositions leads to notable improvements in object removal robustness and background preservation, as discussed and compared in Table 2.
@@ 3.2 ObjectClear
Figure 3: The Framework of ObjectClear. Given an input image and a target object mask, ObjectClear employs an Object-Effect Attention mechanism to guide the model toward foreground removal regions by learning attention masks. The predicted mask further enables an Attention-Guided Fusion strategy during inference, which substantially preserves background details.
Our ObjectClear is built upon SDXL-Inpainting [1]. While SDXL-Inpainting uses a noise map $z_{t}$, a masked image $I_{m}$, a corresponding mask $M_{o}$, and a text prompt $c$ as inputs, we feed the original image $I_{in}$ instead of $I_{m}$ into the model, and our inputs are expressed as a tuple $<z_{t},I_{in},M_{o},c>$. This design encourages the model to better attend to the effect of the target object by leveraging its visual features. Moreover, it facilitates more effective utilization of background information behind the object in cases involving moving transparent objects, such as glass cups. Apart from the input design, we further introduce an Object-Effect Attention Mechanism and an Attention-Guided Fusion Strategy to explicitly attend object-effect regions and enhance background preservation. Detailed descriptions of these components are provided below.
Object-Effect Attention. To enable the model to better attend to both the target object region and its associated visual effect regions, we follow the approach of Li et al. [16] by stacking text and visual object information as guidance for the cross-attention layers in the base model. Specifically, the text prompt is expressed as “remove the instance of”, and the visual object is obtained by applying a dot product between the input image $I_{in}$ and $M_{o}$, as illustrated in Fig. 3(a). These two modalities are then encoded into text embeddings and visual embeddings using the CLIP [25] text and vision encoders, respectively. To unify their representation spaces, the visual embeddings are further projected into the same dimensional space as the text embeddings using a lightweight trainable MLP composed of two linear layers. The resulting text embeddings and projected object embeddings are then stacked and used as guidance in the cross-attention blocks of the base model. To encourage the model to focus more accurately on the object and its effect regions, we introduce a mask loss $\mathcal{L}_{mask}$. Concretely, we extract the cross-attention maps corresponding to the visual embedding tokens and denote them as $\mathbf{A}$, which we supervise with the annotated foreground object-effect masks $M_{fg}$ from our OBER dataset. Following [35, 23], $\mathcal{L}_{\text{mask}}$ is designed to minimize the attention values in the background regions while maximizing those in the foreground. This objective can be formulated as:
@@
where []delimited-[][$\cdot$][ ⋅ ] denotes indexing operator, and $\mathbf{A}[M_{fg}]$ refers to the attention values within the foreground regions specified by the mask $M_{fg}$.
Attention-Guided Fusion. Interestingly, we observe that applying Object-Effect Attention not only improves ObjectClear’s precision in object removal, but also produces attention maps that accurately capture both the object and its effects. Based on this observation, we propose Attention-Guided Fusion strategy that utilizes attention maps during inference to seamlessly combine the generated result with the original image. Specifically, we extract the first-layer cross-attention map corresponding to the object embedding during inference. This attention map is then upsampled to match the spatial resolution of the original image, forming a soft estimate of the object-effect region. To enhance the spatial smoothness and avoid hard edges, we apply a Gaussian blur to the upsampled attention map, producing a softened object-effect mask. This blurred mask is then used to blend the generated image with the original input via a copy-and-paste operation, where pixels within the object-effect region are taken from the generated result, and the rest are preserved from the original image. This strategy significantly reduces undesired background changes caused by the diffusion process and VAE reconstruction, while preserving high-fidelity object removal. Unlike BrushNet [12], which relies on user-provided object masks to guide blending, our approach adaptively produces the object-effect mask from our Object-Affect Attention.

### v2

# Precise Object and Effect Removal with Adaptive Target-Aware Attention

@@ Abstract
Object removal requires eliminating not only the target object but also its associated visual effects such as shadows and reflections. However, diffusion-based inpainting and removal methods often introduce artifacts, hallucinate contents, alter background, and struggle to remove object effects accurately. To address these challenges, we propose ObjectClear, a novel framework that decouples foreground removal from background reconstruction via an adaptive target-aware attention mechanism. This design empowers the model to precisely localize and remove both objects and their effects while maintaining high background fidelity. Moreover, the learned attention maps are leveraged for an attention-guided fusion strategy during inference, further enhancing visual consistency. To facilitate the training and evaluation, we construct OBER, a large-scale dataset for OBject-Effect Removal, which provides paired images with and without object-effects, along with precise masks for both objects and their effects. The dataset comprises high-quality captured and simulated data, covering diverse objects, effects, and complex multi-object scenes. Extensive experiments demonstrate that ObjectClear outperforms prior methods, achieving superior object-effect removal quality and background fidelity, especially in challenging scenarios.
Figure 1: Object Removal Comparison. Given an object mask, prior methods often leave residual artifacts or hallucinate undesirable content, change background, and usually fail to remove associated effects such as shadows and reflections. In contrast, our ObjectClear precisely eliminates both the object and its associated effects, achieving seamless object removal results even in challenging cases.
footnotetext: Equal contribution  Corresponding author
@@ 1 Introduction
Recent advances in generative models [30, 26, 23], have demonstrated strong capabilities in image editing and are widely adopted in real-world applications. Among these tasks, object removal has become a key application, enabling users to erase unwanted content from images. However, seamlessly removing both the object and its associated visual effects (e.g., shadows and reflections) while preserving the background remains a challenging problem, as it requires precisely localizing the object together with its effects, restoring the removed regions with background visual consistency, and maintaining the rest of the image unchanged.
With the rapid development of diffusion-based image generation models [8, 26, 14, 15] and object segmentation techniques [13, 28], combining diffusion-based image generators with target object masks has become the mainstream paradigm for object removal [26, 48, 12, 18, 34, 6, 10, 4]. However, as shown in Figs. 1 and 5, existing methods often fail to remove the visual effects of target objects, generate undesired objects in the removed areas, or unintentionally modify the background (see Fig. 7). These limitations mainly stem from the lack of explicit modeling of the correlation between the target object and its associated effects, as well as the absence of effective constraints to guide the generative model’s attention toward the removal regions.
To address these challenges, we propose ObjectClear, a novel framework designed to achieve precise and complete removal of both objects and their visual effects while preserving background fidelity. To enhance the modeling of object–effect regions, we introduce an Adaptive Target-Aware Attention (ATA) mechanism that learns to localize objects and their associated effects under the supervision of object–effect masks. By integrating text–image embeddings, ATA effectively decouples foreground removal from background reconstruction, enabling the model to remove targets accurately and suppress undesired object generation within the removed regions. During inference, we further introduce an Attention-Guided Fusion (AGF) module that leverages ATA-predicted object–effect attention maps to perform adaptive input–output fusion, preserving fine-grained background details including color and texture. In addition, we introduce a Spatially-Varying Denoising Strength (SVDS) strategy to address incomplete removal and color inconsistency caused by uniform denoising strength. Together, these designs empower ObjectClear to deliver precise and high-fidelity effect-object removal.
@@
To train ObjectClear and fully realize its potential, a dedicated dataset containing paired images with and without objects, along with precise masks for both objects and their effects, is essential. Unfortunately, to the best of our knowledge, no such large-scale training dataset is publicly available. Current object removal datasets can be categorized into simulated and camera-captured data. (1) Simulated data. These datasets are often constructed by copy-pasting objects [11, 17] or using pretrained inpainting models to generate pseudo ground truth [35]. While this allows low-cost generation of large-scale data, such datasets typically lack object effects such as shadows and reflections, causing models trained on them to struggle with effect removal. (2) Camera-captured data. Some works [31, 37] extract paired frames from fixed-viewpoint videos, but this restricts the foreground to moving objects only and makes it difficult to maintain background consistency between paired samples. Others [38, 42, 39] capture image pairs before and after object removal, but these datasets remain small in scale due to their high collection cost and are not publicly released.
@@
To overcome these limitations, we present OBER, a hybrid dataset for OBject-Effect Removal, which combines both camera-captured and simulated data with diverse foreground objects, background scenes, and object effects (e.g., indoor/outdoor shadows and reflections). For the camera-captured data, we annotate precise object and object-effect masks, which serve as critical supervision during training. For the simulated data, we leverage these masks to compute accurate alpha maps, enabling realistic alpha blending of RGBA object layers with effects and high-quality backgrounds. We further extend the simulation pipeline to include multi-object scenarios, enhancing model robustness in challenging cases involving occlusion and object interactions, as shown in the last two rows of Fig. 5. By combining the realism of captured data with the scalability of simulation, OBER offers a high-quality, large-scale, and diverse dataset with a total of 12,715 training samples. In addition, we introduce two new benchmarks: OBER-Test and OBER-Wild, to support future research.
Our contributions are summarized as follows: (1) We introduce a novel framework, ObjectClear, which incorporates an Adaptive Target-Aware Attention (ATA) mechanism that adaptively focuses on foreground removal regions, together with an Attention-Guided Fusion (AGF) and a Spatially-Varying Denoising Strength (SVDS) strategy, thereby improving object removal quality and background fidelity. (2) We propose OBER, a high-quality and large-scale hybrid dataset for object-effect removal, featuring diverse objects, fine-grained annotations of object-effect masks, and complex scenarios across both simulated and camera-captured settings. (3) ObjectClear achieves superior performance on both synthetic and real-world benchmarks, demonstrating clear advantages in quantitative results and visual quality.
@@ 2 Related Work
Image Inpainting. Image inpainting is a long-standing visual editing task that aims to seamlessly reconstruct pixels within a given mask. Early approaches predominantly adopt generative adversarial networks (GANs) [7], but often suffer from limited realism and diversity [19, 25, 29, 43]. With the rapid advances in diffusion models [30, 33, 9, 26], many methods [1, 22, 23, 48, 12, 41, 36] have begun leveraging their strong generative priors to synthesize high-fidelity content, achieving state-of-the-art results in image inpainting. In this work, we adapt the SDXL-Inpainting model [26] for photorealistic completion. However, despite their strong generative capabilities, existing inpainting models often lack awareness of object effects (e.g., shadows and reflections), leading to incomplete or inconsistent object removal results.
Object Removal. Object removal is a specialized branch of image inpainting that often requires explicit consideration of object effects (e.g., shadows and reflections) to achieve complete removal, a problem increasingly dominated by diffusion-based methods [38, 11, 21, 42, 18, 3, 48, 6, 4, 39, 10, 34, 37, 47], and recently extended to the video domain [46, 16, 24, 32]. A common strategy is to curate triplet datasets. ObjectDrop [38] builds a real-world dataset by capturing the same scene before and after object removal, but its limited scale and non-public availability hinder broader adoption. To improve data scalability, methods like SmartEraser [11] and EraseDiffusion [21] rely on synthetic datasets generated by segmentation or matting methods to extract foreground objects, but these datasets typically lack annotations of object effects, limiting the ability to remove shadows or reflections. OmniPaint [42] further expands its dataset by automatically annotating unlabeled images using a model trained on small-scale real data, while RORem [18] also incorporates human annotators to ensure annotation quality. In parallel, works like RORD [31] and OmniEraser [37] scale data generation by mining realistic video frames with fixed viewpoints, selectively pairing frames w/ and w/o target objects while preserving natural object effects. GeoRemover [47] leverages depth geometry to achieve precise object removal. To eliminate reliance on curated datasets altogether, other methods opt for test-time optimization [4, 34, 10] or providing plug-and-play solutions [6]. However, these models rely on implicitly learning object effects without explicitly modeling effect maps, which makes it difficult to maintain background consistency.
Figure 3: The Framework of ObjectClear. Given an input image and a target object mask, ObjectClear employs an Adaptive Target-Aware Attention mechanism to guide the model toward foreground removal regions by learning attention masks. The predicted mask further enables an Attention-Guided Fusion strategy during inference, which substantially preserves background details.
@@ 3 Methodology
To achieve accurate removal of target objects and their associated effects, we propose a comprehensive framework with a data curation pipeline (OBER Dataset, Sec. 3.1) and a dedicated object removal model (ObjectClear, Sec. 3.2).
@@ 3.1 OBER Dataset
To learn an object–effect removal model, a high-quality dataset with paired images and precise masks for both objects and their effects is required. However, to our knowledge, no such large-scale dataset is publicly available. To this end, we introduce the OBER dataset (OBject-Effect Removal), designed to balance data realism and scalability. As shown in Fig. 2, OBER is a hybrid dataset consisting of two parts: (1) a small set (2,878) of camera-captured images adhering to physical realism, and (2) a larger set (10,000) of simulated images generated by compositing foreground objects, extracted from (1), onto diverse background scenes.
Camera-Captured Data. (1) Capture Paired Images. Following the approach of ObjectDrop [38], we use fixed cameras to construct a counterfactual dataset by capturing each scene before and after the removal of a single object, while keeping all the other factors unchanged. For each pair, the image with the object is used as the input $I_{in}$, and the image without the object serves as the ground truth $I_{gt}$. In total, we collected 2,878 such counterfactual pairs with 2,715 for training and 163 for testing. Our dataset encompasses a wide variety of everyday objects commonly found in indoor (e.g., benches, cups, dolls) and outdoor scenes (e.g., pedestrians, vehicles). These image pairs also preserve visual effects such as shadows and reflections. To reduce potential pixel-level misalignment, we downsample and crop all images to a fixed training resolution (i.e., 512 $\times$ 512). (2) Segment Object Mask $M_{o}$. To obtain the object mask $M_{o}$ as network input alongside the input image $I_{in}$, we apply off-the-shelf detection and segmentation models, such as DINO [44] and SAM [13] to $I_{in}$. (3) Segment Object-Effect Mask $M_{o}\cup M_{e}$. Different from previous methods, we propose to explicitly model the object effects. Therefore, we introduce an object-effect mask $M_{fg}$ that covers both the object $M_{o}$ and its associated visual effects $M_{e}$ in our dataset. Unlike the coarse object-effect masks provided in the RORD dataset [31], we construct $M_{fg}$ in an efficient and accurate way by computing the pixel-wise difference between $I_{in}$ and $I_{gt}$. Pixels with differences above a predefined threshold are regarded as part of the object-effect mask. This object-effect mask provides crucial supervision during training, allowing the network to adaptively learn to focus on the target removal regions, including both the object itself and its associated effects.
@@
$\alpha(p)=\begin{cases}0,&\text{if }p\in\text{$\overline{M_{o}}\cap\overline{M_{e}}$}\\ 1,&\text{if }p\in\text{$M_{o}$}\\ (I_{\text{gt}}-I_{\text{in}})/(I_{\text{gt}}+\varepsilon),&\text{if }p\in\text{$M_{e}$}\\ \end{cases},$ (1)
where $M_{o}$ and $M_{e}$ are object and effect regions, respectively.
(3) Blend Objects with Backgrounds. We randomly sample a background image $I_{bg}$ and a foreground object image with effects $I_{oe}$, and synthesize a composite image using alpha blending: $I_{comp}=(1-\alpha)\cdot I_{bg}+\alpha\cdot I_{oe}$. Beyond single-object cases, we also synthesize multi-object data by compositing multiple foreground objects with the same lighting directions, thereby covering scenarios involving object occlusions, as illustrated in Fig. 2(b). These designs ensure physically plausible placement and consistent lighting, reducing the domain gap between simulated and real data. Synthetic data also overcomes challenges in real-world collection, such as limited scale and complex cases (e.g., multi-object occlusions, reflections), while its controllability expands coverage and diversity. In total, we generate 10,000 composite images, significantly enriching the diversity and scalability of our dataset. In particular, the simulation of multi-object compositions leads to notable improvements in object removal robustness and background preservation, as discussed and compared in Table 2.
@@ 3.2 ObjectClear
Our ObjectClear is built upon SDXL-Inpainting [26]. While SDXL-Inpainting uses a noise map $z_{t}$, a masked image $I_{m}$, a corresponding mask $M_{o}$, and a text prompt $c$ as inputs, we feed the original image $I_{in}$ instead of $I_{m}$ into the model, and our inputs are expressed as a tuple $z_{t},I_{in},M_{o},c>$. This design encourages the model to better attend to the effect of the target object by leveraging its visual features. Moreover, it facilitates more effective utilization of background information behind the object when transparent objects are to be removed, such as glass cups. To achieve precise and complete object removal while enhancing background preservation, we introduce Adaptive Target-Aware Attention (ATA), Attention-Guided Fusion (AGF), and Spatially-Varying Denoising Strength (SVDS) strategy, which explicitly attend to object-effect regions.
Adaptive Target-Aware Attention. To enable the model to better attend to both the target object region and its associated effect regions, we integrate text and object image as an object prompt for the cross-attention layers in the base model. Specifically, the text prompt is expressed as “remove the instance of”, and the visual object is obtained by applying a dot product between the input image $I_{in}$ and $M_{o}$, as illustrated in Fig. 3(a). These two modalities are then encoded into text embeddings and visual embeddings using the CLIP [27] text and vision encoders, respectively. To unify their representation spaces, the visual embeddings are further projected into the same dimensional space as the text embeddings using a lightweight trainable MLP composed of two linear layers. The resulting text embeddings and projected object embeddings are then stacked and used as guidance in the cross-attention blocks of the base model. To encourage the model to focus more accurately on the object and its effect regions, we introduce a mask loss $\mathcal{L}_{mask}$. Concretely, we extract the cross-attention maps corresponding to the visual embedding tokens and denote them as $\mathbf{A}$, which we supervise with the annotated foreground object-effect masks $M_{fg}$ from our OBER dataset. $\mathcal{L}_{\text{mask}}$ is designed to minimize the attention values in the background regions while maximizing those in foreground. This objective can be formulated as:
@@
where $\cdot$] denotes indexing operator, and $\mathbf{A}[M_{fg}]$ refers to the attention values within the foreground regions specified by the mask $M_{fg}$.
Attention-Guided Fusion. Interestingly, we observe that applying Adaptive Target-Aware Attention not only improves ObjectClear’s precision in object removal, but also produces attention maps that accurately capture both the object and its effects. Based on this observation, we propose an Attention-Guided Fusion strategy that leverages the predicted attention maps during inference to seamlessly blend the generated result with the original image. Specifically, we extract the first-layer cross-attention map corresponding to the object embedding during inference. This attention map is then upsampled to match the resolution of the original image, forming a soft estimate of the object-effect region. To avoid edge artifacts during blending, we apply a Gaussian blur to the upsampled attention map, producing a soft-edged object-effect mask. This mask is then used in an alpha-blending operation to fuse the generated image with the original input. This strategy significantly reduces undesired background changes introduced by the diffusion denoising process and VAE reconstruction, achieving high-fidelity object removal. Unlike BrushNet [12], which relies on user-provided object masks that usually exclude effect regions, our approach leverages precise object-effect masks adaptively generated by Adaptive Target-Aware Attention.
Spatially-Varying Denoising Strength. In diffusion-based image editing, the initial latent for denoising is typically obtained by adding a certain amount of noise to the input image latent. The denoising strength (DS) $\mathrm{DS}\in[0,1]$ controls the noise level: a larger value injects more noise, thereby pushing the initial latent $z_{t}$ closer to the pure-noise prior. Specifically, when $\mathrm{DS}=1.0$, the diffusion process starts entirely from noise and discards all information from the inputs, achieving complete object removal but often introducing noticeable global color shifts. In contrast, using a slightly lower denoising strength (e.g., $\mathrm{DS}=0.99$) preserves color consistency but risks incomplete object removal or hallucinated unwanted objects. Motivated by these observations, we propose Spatially-Varying Denoising Strength (SVDS), which applies $\mathrm{DS}=1.0$ within the masked object region and $\mathrm{DS}=0.99$ in the unmasked background, via re-injecting the background during inference. This strategy achieves complete object-effect removal while maintaining background color consistency and preventing edge artifacts during Attention-Guided Fusion (see ablation in the supp.).